A sphere is to be painted. How much surface area must be covered if the radius is 3.6 feet? Leave the answer in terms of pie.
Answer by stanbon(75887)

A sphere is to be painted. How much surface area must be covered if the radius is 3.6 feet? Leave the answer in terms of pie.
--------------------
Sarface Area of a sphere = 4(pi)r^2
-----
Your Problem:
SA = 4(pi)3.6^2 = 51.84pi sq.ft.
================
Cheers,
Stan H.